# Quantum-Classical Hybrid Quantized Neural Network

### v2

@@ Results
To verify the advantages of the training on neural network, we undertake the experiment using coherent Ising machines for the classification of coat and sandal images from the Fashion MNIST dataset. The dataset consists of $60,000$ training images and $10,000$ test images, each representing a grayscale clothing item from $10$ classes, including T-shirts/tops, trousers, pullovers, dresses, coats, sandals, shirts, sneakers, bags, and ankle boots. Each image is $28\times 28$ pixels, with pixel values ranging from 00 to $255$. Additional details of the experimental results are deferred to the supplementary.
@@
To evaluate the performance of our quantum-based approach for training quantized neural networks, we present a series of comparative analyses. Figure 3(b) compares the test accuracy and running time of our model and quantum approach against previous classical approaches, including straight through estimator (STE) [47] and BinaryConnect [48]. We provide a brief description of the two classical algorithms in the supplementary. Here the proposed quantum-based method achieves a test accuracy of $94.95\%$ with a significantly reduced running time of $0.185$ ms, demonstrating superior efficiency and effectiveness.
Further, Figure 3(c) examines the memory demand and inference complexity, contrasting a neural network with continuous weights and sigmoid activation functions against our quantized approach, which approximates the activation function with a simple piecewise linear function. Our method reduces memory cost to 2.75 bytes and inference time to 0.0762 seconds while maintaining a test accuracy of $94.95\%$, highlighting its resource efficiency. These results collectively underscore the advantages of leveraging quantum computing for training quantized neural networks, offering improvements in speed, and resource efficiency.
@@
Algorithm Activation FP Model 4-bit 3-bit 2-bit
RTN [49] ReLU 98.7% 95.4% 98.7% 50.0%
LeakyReLU 98.7% 98.7% 98.7% 75.4%
Sigmoid 98.7% 97.2% 95.4% 75.4%
AdaRound [9] ReLU 98.7% 98.7% 98.8% 50.0%
LeakyReLU 98.7% 98.7% 98.7% 98.7%
Sigmoid 98.7% 97.2% 95.4% 50.0%
DFQ [50] ReLU 98.7% 95.4% 98.7% 50.0%
LeakyReLU 98.7% 98.7% 98.7% 98.7%
Sigmoid 98.7% 98.7% 98.7% 50.0%
Table 2: Classfication accuracy (%) under different PTQ algorithms, activation functions, and bit widths.
While most methods maintain high accuracy at $4$-bit and $3$-bit precision, a significant drop is observed at $2$-bit, especially with ReLU activations. Specifically, our method achieves $94.95\%$ test accuracy with an effective parameter precision of just $1.1$ bits, substantially outperforming almost all listed $2$-bit baselines. This result highlights the potential of our framework under aggressive quantization.
Spline interpolation requires careful consideration of the number of intervals, as it directly influences qubit resource allocation and must be optimized for quantum hardware constraints. In our experimental setup, we partitioned the input range into four intervals, achieving an optimal trade-off between qubit efficiency and segmentation accuracy. The corresponding breakpoints, defined at $[-8,-c,0,c,8]$, yielded the following results:
@@
Figure 4 illustrates the convergence behavior of the Quantum Conditional Gradient Descent algorithm when applied to the neural network training utilizing a hybrid quantum-classical approach with CIM and Gurobi solvers. The algorithm achieves convergence to the optimal solution within 321 iterations, with the objective residual and constraint residual presented on a logarithmic scale. The CIM solver exhibits consistend convergence behavior, steadily driving both the objective and constraint residual below the threshold value. In contrast, the Gurobi solver maintains convergence but displays oscillatory patterns in the constraint residual between iterations 30-100. Notably, the hybrid strategy achieves convergence in just 321 iterations, demonstrating its computational efficiency by effectively combining the strengths of CIM and the classical approach. To further validate the robustness and reproducibility of these results, we performed five independent experiments achieving convergence to the optimal solution in 723, 338, 263, 432, and 427 iterations, respectively. These consistent outcomes across multiple trials confirms its reliability and reproducibility in practical applications.
Robustness of hybrid approach: impact of coefficient matrix rounding
To investigate the convergence behavior of the hybrid algorithm in neural network training under different matrix rounding precision levels, we systematically rounded the matrix coefficients to digit precisions ranging from 1 to 10 decimal places. The algorithm was then executed for each precision setting until convergence to the optimal solution was achieved. Specifically, matrix elements were rounded to 1 through 10 significant digits, and the corresponding number of iterations required to attain the optimal solution at each precision level was recorded as follows:
These results indicate the variability in convergence speed, with lower precision levels (e.g., 1-3 digits) requiring substantially more iterations (up to 43,343 for 3 digits), likely due to increased numerical instability and larger initial differences. In contrast, higher precision levels (e.g., 5-10 digits) converge more efficiently, with the minimum iterations observed at 2904 for 10 digits, suggesting that increased precision enhances the convergence stability.
Precision (Digits) 1 2 3 4 5 6 7 8 9 10
Iterations to optimal 6303 18130 43343 4746 3993 3787 6099 6643 5871 2904
Table 4: Convergence speed of QCGD algorithm across precision levels
Figure 5(a) displays the QUBO matrix evolution using circular heatmaps, highlighting the impact of precision levels on iterative trajectories. Despite the differences revealed in these trajectories, our results confirm that all trajectories converge to the optimal solution, consistent with the theoretical convergence of the QCGD algorithm. Notably, even under minimal precision conditions (e.g., 1-digit rounding), matrix adjustments in later iterations remain marginal. These consistently small perturbations across all precision levels suggest that a lazy evaluation strategy for QCGD (Quantum-Constrained Gradient Descent) could be viable. Such an approach would minimize redundant quantum circuit executions, thereby improving computational efficiency on quantum hardware.
The smoothed absolute differences of the objective value and constraint residual relative to the 10-digit precision baseline were recorded, with results visualized in Figure 5(b) and 5(c). This setup allowed us to assess the impact of numerical precision on convergence trajectories. Both subplots demonstrate that differences across all precision levels remain small, typically less than $10^{-2}$, indicating the hybrid algorithm’s robustness to precision variations. A downward trend is observed in the difference of constraint residual, with differences decreasing sharply after initial fluctuations and stabilizing post-1500 iterations, confirming convergence to a consistent optimal solution regardless of precision, as supported by prior results. The rapid initial decline in the constraint residual difference, particularly for lower precisions (e.g., 1-3 digits), suggests early sensitivity to numerical accuracy, while the differences remain stably small in later iterations. The near-overlap of higher precision curves (e.g., 7-9 digits) highlights diminishing returns from increased precision.
@@ Methods
The QCGD method is build on copositive programming (CP) [51] which is a mathematical optimization branch that has garnered significant attention for its applications in various optimization problems. In copositive programming, the objective is to optimize over the cone of copositive matrices, which are matrices for which the quadratic form yields nonnegative values over the nonnegative orthant. One key application of copositive programming is in transforming binary quadratic problems into completely positive problems. This transformation involves reformulating the standard quadratic problem as a CP program, enabling the optimization of a quadratic form over the Cartesian product of simplices.
@@
Consider the following quadratic 00-$1$ programming problem with quadratic constraints programming,
@@
$\mathbf{x}^{T}\mathbf{Q}_{t}\mathbf{x}\leq c_{t},1\leq t\leq T$ (37)
@@
$\sum_{\ell=1}^{\bar{\ell}}2^{\ell}\cdot\delta^{(\ell)}_{a^{l}_{x,k}}+(2^{\bar{\ell}+1}-2)\cdot w^{l}_{jk}-z^{l}_{x,j,k}\leq 2^{\bar{\ell}+1}-2,$ (38)
$z^{l}_{x,j,k}\leq\sum_{\ell=1}^{\bar{\ell}}2^{\ell}\cdot\delta^{(\ell)}_{a^{l}_{x,k}},$ (39)
$z^{l}_{x,j,k}\leq(2^{\bar{\ell}+1}-2)\cdot w^{l}_{jk}.$ (40)
@@
    where $\mathbf{w}^{(k)}$ is the $k$-th solution. Set the update direction as $\mathbf{H}_{t}=\mathbf{w}_{t}\mathbf{w}_{t}^{\top}$.
@@
    Update the solution: $\mathbf{V}_{t+1}=(1-\eta_{t})\mathbf{V}_{t}+\eta_{t}\mathbf{H}_{t}$.
@@
For problem
$\min_{\mathbf{D}}\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}\mathbf{D})$ (42)
over the cone of completely positive matrices, if $\mathbf{D}_{t}$ obtained from the oracle satisfies the following condition:
$\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}(\mathbf{D}_{t}-\mathbf{V}_{t}))\leq\delta\cdot\mathrm{Tr}(\mathbf{Q}^{(t)}_{\text{QUBO}}(\mathbf{D}^{*}_{t}-\mathbf{V}_{t}))+\frac{\xi_{t}}{\sqrt{t}},$ (43)
where $\mathbf{D}^{*}_{t}$ is the exact optimal solution, $\delta\in(0,1]$ denotes relative error bound, $\xi_{t}$ represents the random additive error with $\mathbb{E}[\xi_{t}]\leq\varepsilon$. Then $\mathbf{D}_{t}$ is said to be an output of a $(\delta,\varepsilon)$-Inexact oracle.
A quantum computer without any imperfections corresponds to the scenario where $\delta=1$ and $\varepsilon=0$. The following lemma indicates that convergence can still be achieved even if the single quantum computation acts as an inexact oracle. This theoretical framework ensures that the overall algorithm remains robust despite potential inaccuracies in the quantum computation. Theoretical proof follows from that in [45], and we include it in the supplementary for completeness.
@@
[…]
@@
$||\mathbf{Q}^{(t+1)}_{\text{QUBO}}-\mathbf{Q}^{(t)}_{\text{QUBO}}||_{F}<\Delta_{t},$ (46)
@@ Conclusion
In conclusion, this paper introduces a QUBO-based framework that enables arbitrary activation and loss functions through spline interpolation, significantly expanding the applicability of quantum computing in neural network training. The theoretically derived upper bound on Ising spin requirements, along with the empirically validated convergence of our hybrid algorithm, even under randomized outputs and limited coefficient matrix precision, further validates the robustness of our approach. The comparison with classical methods highlights the superior performance of the Ising machine in quantized neural network training, while the hybrid algorithm demonstrates consistent efficacy and stability. Meanwhile, our model offers promising potential for applications demanding trainable activation functions, paving the way for quantum-enhanced neural architectures. This work represents a pivotal advancement toward harnessing the full potential of quantum-accelerated machine learning in real-world implementations.
@@
Appendix A Error Bound of Piecewise Linear Approximation in FNN

### v4

@@ Results
To verify the advantages of the training on neural networks, we undertake the experiment using coherent Ising machines for the classification of coat and sandal images from the Fashion MNIST dataset. The dataset consists of $60,000$ training images and $10,000$ test images, each representing a grayscale clothing item from $10$ classes, including T-shirts/tops, trousers, pullovers, dresses, coats, sandals, shirts, sneakers, bags, and ankle boots. Each image is $28\times 28$ pixels, with pixel values ranging from 0 to $255$. Additional details of the experimental results are deferred to Section 6.1 of the supplementary.
@@
To evaluate the performance of our quantum-based approach for training quantized neural networks, we present a series of comparative analyses. Figure 4(b) compares the test accuracy and running time of our model and quantum approach against previous classical approaches, including straight through estimator (STE) [56] and BinaryConnect [57]. Here, the proposed quantum-based method achieves a test accuracy of $94.95\%$ with a significantly reduced running time of $0.185$ ms, demonstrating superior efficiency and effectiveness.
@@
Further, Figure 4(c) examines the memory demand and inference complexity, contrasting a neural network with continuous weights and sigmoid activation functions against our quantized approach, which approximates the activation function with a simple piecewise linear function. Our method reduces memory cost to 2.75 bytes and inference time to 0.0762 seconds while maintaining a test accuracy of $94.95\%$, highlighting its resource efficiency. These results collectively underscore the advantages of leveraging quantum computing for training quantized neural networks, offering speed improvements and resource efficiency.
During inference with piecewise linear activation, the activation value at each neuron is determined in two steps: first by computing the pre-activation input, and then by identifying the interval within which this input falls. This interval-indexing step, or equivalently, locating the unique $\beta_{i}=1$, allows the model to evaluate the activation function through the corresponding piecewise linear segment. This process is computationally more efficient than evaluating a smooth nonlinear activation function (e.g., sigmoid or $\tanh$), which typically involves multiple floating-point operations such as exponentiation, division, and additional multiplications arising from polynomial or rational approximations used in hardware implementations.
In contrast, once the correct interval is determined, a piecewise linear activation requires only one interval lookup, one multiplication, and one addition. The time per activation evaluation can therefore be expressed as $T_{\mathrm{PWL}}=T_{\mathrm{lookup}}+T_{\mathrm{mul}}+T_{\mathrm{add}},$ which is lower than the cost of evaluating a smooth nonlinear activation, such as a sigmoid in the form $T_{\mathrm{sigmoid}}=T_{\exp}+T_{\mathrm{add}}+T_{\mathrm{div}}$. Consequently, spline-based piecewise linear activations not only facilitate a QCBO formulation during training but also offer a tangible acceleration during inference.
@@
Algorithm Activation FP Model 4-bit 3-bit 2-bit 1-bit
RTN [58] ReLU 98.8% 97.73% 97.73% 73.35% 50.00%
LeakyReLU 98.8% 98.70% 98.62% 76.77% 50.00%
Sigmoid 98.8% 98.76% 98.34% 55.84% 50.00%
AdaRound [9] ReLU 98.8% 97.73% 97.76% 69.87% 50.00%
LeakyReLU 98.8% 98.68% 98.78% 71.86% 50.00%
Sigmoid 98.8% 98.77% 98.61% 66.60% 50.00%
DFQ [59] ReLU 98.8% 96.62% 96.07% 84.99% 64.10%
LeakyReLU 98.8% 98.08% 97.80% 88.50% 61.40%
Sigmoid 98.8% 98.75% 96.44% 60.80% 50.00%
GPTQ [60] ReLU 98.8% 97.65% 97.50% 92.84% 50.00%
LeakyReLU 98.8% 98.80% 98.11% 91.82% 50.01%
Sigmoid 98.8% 98.80% 98.64% 91.21% 50.00%
ZeroQuant [61] ReLU 98.8% 97.79% 97.24% 93.59% 82.50%
LeakyReLU 98.8% 98.80% 98.10% 96.23% 91.43%
Sigmoid 98.8% 98.80% 97.28% 91.97% 91.45%
SubsetQ [62] ReLU 98.8% 97.79% 97.78% 96.70% 88.90%
LeakyReLU 98.8% 98.80% 98.65% 95.94% 91.45%
Sigmoid 98.8% 98.80% 98.80% 94.98% 91.45%
Table 2: Classification accuracy (%) under different PTQ algorithms, activation functions, and bit widths.
Across all post-training quantization baselines evaluated, including RTN [58], AdaRound [9], DFQ [59], GPTQ [60], ZeroQuant [61], and SubsetQ [62], accuracy degrades as the bit-width approaches the extremely low-precision regime ($\leq$2 bits). Conventional 1-bit and 2-bit post-training quantization methods often lead to severe representational collapse, with most failing to surpass $90\%$ accuracy and many performing no better than random chance. Against this backdrop, our quantum-trained model maintains desirable decision boundaries even at 1.1‑bit precision, attaining an accuracy of $94.95\%$.
In the low bit width regime, SubsetQ delivers competitive performance. However, it suffers from several inherent limitations. Notably, its computational cost grows sharply with increasing bit-width, because the method must evaluate across a candidate quantization set whose size scales as $2^{b}$. This exponential dependence renders SubsetQ efficient only in low precision regime (e.g., 121–2 bits), and increasingly prohibitive at higher bit‑widths. In contrast, our quantum‑optimization approach does not rely on enumerating quantization candidates and scales more favorably with target precision, thereby enabling flexible training at arbitrary effective bit‑widths, including the 1.1‑bit configuration employed in our experiments.
Spline interpolation requires careful consideration of the number of intervals, as it directly influences qubit resource allocation and must be optimized for quantum hardware constraints. In our experimental setup, we partitioned the input range into four intervals, achieving an optimal trade-off between qubit efficiency and segmentation accuracy. The corresponding breakpoints, defined at [8,c,0,c,8][-8,-c,0,c,8], yielded the following results:
@@
Figure 5 illustrates the convergence behavior of the Quantum Conditional Gradient Descent algorithm when applied to the neural network training utilizing a hybrid quantum-classical approach with CIM and Gurobi solvers. The algorithm achieves convergence to the optimal solution within 321 iterations, with the objective residual and constraint residual presented on a logarithmic scale. To further validate the robustness and reproducibility of these results, we performed five independent experiments, achieving convergence to the optimal solution in 723, 338, 263, 432, and 427 iterations, respectively. These consistent outcomes across multiple trials confirm its reliability and reproducibility in practical applications. In Section 6.2 of the supplementary, we also investigate the impact of coefficient matrix rounding and robustness of hybrid approach.
Effectiveness of DLBO framework
After solving the DLBO relaxation, we obtain a fractional parameter vector $\bar{x}\in[0,1]^{n}$. To obtain a quantized network, in this experiment, we apply an independent randomized Bernoulli rounding, i.e., for each coordinate $j$, we draw $X_{j}\sim\mathrm{Bernoulli}(\bar{x}_{j})$, and interpret $X_{j}\in\{0,1\}$ as the final value of the $j$-th parameter bit. In practice, we repeat this randomized rounding procedure $K$ times, resulting in $K$ candidate binary networks $X^{(1)},\dots,X^{(K)}$. For each candidate, we evaluate the recourse objective $F(X^{(k)})$, i.e., the sum of per–sample losses measuring the discrepancy between the predicted labels and the true labels), and we report the best value as $F_{\min}:=\min_{1\leq k\leq K}F\bigl(X^{(k)}\bigr)$. Theoretical analysis in Section 6.3 of the supplementary shows that this strategy yields a discrete solution whose objective value concentrates around the DLBO relaxation value $F_{\mathrm{DLBO}}$ and improves with increasing $K$ with high probability.
The predictive performance of the DLBO is summarized in Table 4. Across all three datasets, Fashion MNIST [63], Wine [64], and Digits [65], the discrete networks obtained via randomized rounding maintain high approximation ratios, indicating that the performance of the learned binary structures remains extremely close to the relaxed optimum. Correspondingly, the classification performance remains competitive, with accuracies in the range of $90.24\%–96.30\%$ and $90.15\%–94.94\%$ on the test and full datasets, respectively. Weighted precision, recall, and F1 scores exhibit similar behavior, all falling within a relatively high numerical range, demonstrating that the DLBO framework produces stable and well-behaved decision boundaries after rounding.
Table 4: Performance Comparison on Test and Full Datasets
Test Set Full Set
Fashion MNIST Wine Digits Fashion MNIST Wine Digits
Approximation Ratio $0.9414$ $1.0000$ $1.0000$ $0.9414$ $1.0000$ $1.0000$
Accuracy $0.9024$ $0.9167$ $0.9630$ $0.9015$ $0.9494$ $0.9460$
Weight Avg. precision $0.9025$ $0.9235$ $0.9641$ $0.9015$ $0.9518$ $0.9519$
Weight Avg. recall $0.9024$ $0.9167$ $0.9630$ $0.9015$ $0.9494$ $0.9460$
Weight Avg. F1 Score $0.9024$ $0.9168$ $0.9630$ $0.9015$ $0.9496$ $0.9458$
@@ Methods
The QCGD method is built on copositive programming (CP) [47], which is a mathematical optimization branch that has garnered significant attention for its applications in various optimization problems. In copositive programming, the objective is to optimize over the cone of copositive matrices, which are matrices for which the quadratic form yields nonnegative values over the nonnegative orthant. One key application of copositive programming is in transforming binary quadratic problems into completely positive problems. This transformation involves reformulating the standard quadratic problem as a CP program, enabling the optimization of a quadratic form over the Cartesian product of simplices.
@@
Consider the following quadratic 0-11 programming problem with quadratic constraints,
@@
$\mathbf{x}^{T}\mathbf{Q}_{t}\mathbf{x}\leq c_{t},1\leq t\leq T.$ (75)
@@
$\sum_{\ell=1}^{\bar{\ell}}2^{\ell}\,\delta^{(\ell)}_{a^{l}_{x,k}}+(2^{\bar{\ell}+1}-2)w^{l}_{jk}-z^{l}_{x,j,k}$ $\leq 2^{\bar{\ell}+1}-2,\quad\forall x,j,k,$ (76)
$z^{l}_{x,j,k}$ $\leq\sum_{\ell=1}^{\bar{\ell}}2^{\ell}\,\delta^{(\ell)}_{a^{l}_{x,k}},\quad\forall x,j,k,$ (77)
$z^{l}_{x,j,k}$ $\leq(2^{\bar{\ell}+1}-2)w^{l}_{jk},\quad\forall x,j,k.$ (78)
@@
    where $\mathbf{w}^{(k)}$ is the $k$-th solution. Set the update direction as $\mathbf{D}_{t}=\mathbf{w}_{t}\mathbf{w}_{t}^{\top}$.
@@
    Update the solution: $\mathbf{V}_{t+1}=(1-\eta_{t})\mathbf{V}_{t}+\eta_{t}\mathbf{D}_{t}$.
@@
Consider the problem
$\min_{\mathbf{V}}\ \mathrm{Tr}\!\left(\mathbf{Q}^{(t)}_{\mathrm{QUBO}}\mathbf{V}\right)$ (80)
over the cone of completely positive matrices. An oracle is called a $(\delta,\varepsilon)$-inexact oracle if at iteration $t$ it outputs a matrix $\mathbf{D}_{t}$ such that
$\mathrm{Tr}\!\left(\mathbf{Q}^{(t)}_{\mathrm{QUBO}}(\mathbf{D}_{t}-\mathbf{V}_{t})\right)\;\leq\;\delta\cdot\mathrm{Tr}\!\left(\mathbf{Q}^{(t)}_{\mathrm{QUBO}}(\mathbf{D}^{*}_{t}-\mathbf{V}_{t})\right)\;+\;\frac{\xi_{t}}{\sqrt{t}},$ (81)
where $\mathbf{D}^{*}_{t}$ denotes the exact optimal solution of the QUBO problem, $\delta\in(0,1]$ is a relative error parameter, and $\xi_{t}$ is a random additive error satisfying $\mathbb{E}[\xi_{t}]\leq\varepsilon$.
A quantum computer without any imperfections corresponds to the scenario where $\delta=1$ and $\varepsilon=0$. The following lemma indicates that convergence can still be achieved even if the single quantum computation acts as an inexact oracle. This theoretical framework ensures that the overall algorithm remains robust despite potential inaccuracies in the quantum computation. Theoretical proof follows from that in [49], and we include it in Section 5 of the supplementary for completeness.
@@
Here $\text{Objective-gap}_{T}$ denotes the objective value of the solution obtained at the $T$-th iteration and the optimal objective value; $\text{Infeasibility}_{T}$ represents the distance from the solution obtained at the $T$-th iteration to the feasible region. In addition, the bound in (82) holds with probability at least $1-\frac{\max\mathrm{Var}[\xi_{s}]}{\varepsilon^{2}T}$.
@@
[…]
@@ Conclusion
In conclusion, this paper introduces a QUBO-based and convex formulation framework that enables arbitrary activation and loss functions through spline interpolation, significantly expanding the applicability of quantum computing and the understanding of optimization landscape in neural network training. The theoretically derived upper bound on Ising spin requirements, along with the empirically validated convergence of our hybrid algorithm, even under randomized outputs and limited coefficient matrix precision, validates the robustness of our approach. We also establish scalable hybrid training algorithms that decomposes the global QCBO problem into tractable quantum subproblems while preserving the tightness of the solution. Comparisons with classical methods highlights the superior performance of the Ising machine in quantized neural network training, while the hybrid algorithm demonstrates consistent efficacy and stability. Meanwhile, our model offers promising potential for applications demanding trainable activation functions, paving the way for quantum-enhanced neural architectures. This work represents a pivotal advancement toward harnessing the full potential of quantum-accelerated machine learning in real-world implementations.
@@
Appendix A Related Work
In addition to the aforementioned works, other research also explores the intersection of quantum computing and neural networks. Ref. [68] presents a novel training approach based on Adiabatic Quantum Computing (AQC), which utilizes principles of adiabatic evolution to address optimization challenges. The proposed universal AQC scheme is designed for implementation on gate-based quantum computers. A hybrid strategy introduced in Ref. [69] combines quantum and classical methods to accelerate the training of binary neural networks (BNNs). Its quantum component employs the HHL algorithm to solve linear systems of equations for linear regression within a single-layer BNN. The primary contribution of Ref. [70] lies in demonstrating the transfer of a trained artificial neural network to a quantum computing environment. By integrating deep-learned parameters and layer structures, the authors formulate a quadratic binary model suitable for quantum annealing that aligns with the behavior of a classical neural network.
Training quantized neural network.
Neural network quantization techniques can be broadly categorized into quantization-aware training (QAT) and post-training quantization (PTQ). In QAT, quantization effects are simulated during training so that the optimizer learns weights that are intrinsically robust to low-bit inference. Early QAT methods include BinaryConnect [57], which constrains weights to binary values during the forward pass while maintaining full-precision copies for gradient updates. A key ingredient enabling QAT is the straight-through estimator (STE) [56, 71], which approximates gradients through discrete operations by substituting their non-differentiable Jacobians with surrogate identities. While QAT generally achieves superior accuracy, it requires full retraining, extensive data access, and increased computational cost.
In contrast, PTQ methods quantize a pre-trained model without full retraining and often use only a small calibration set. A straightforward baseline is round-to-nearest (RTN), which simply scales weights and rounds each to the nearest fixed-point value [58]. More sophisticated PTQ methods optimize the quantization procedure itself. AdaRound, for example, learns optimal weight rounding offsets by formulating the rounding decision as a local loss minimization (approximated as a QUBO problem) and solving it with a continuous relaxation [9]. Other approaches introduce lightweight fine-tuning steps into PTQ. For instance, layer-wise differentiable calibration methods adjust weights by backpropagating on a small calibration set to minimize the error between each layer’s quantized outputs and its full-precision outputs [72]. The recent GPTQ algorithm pushes this further by using second-order information, which applies an approximate Hessian-based optimizer to choose quantized weight values that incur minimal loss increase [60], enabling PTQ of large transformers down to 343–4 bits with negligible accuracy degradation. Several PTQ techniques target transformer models specifically. ZeroQuant combines dynamic, per-token activation quantization to suppress outliers with fine-grained group-wise weight quantization to maintain model accuracy on language tasks [61]. Meanwhile, SubsetQ introduces a non-uniform quantizer that selects an optimal subset of quantization levels from a larger universal set, allowing flexible step sizes tuned to each layer’s distribution [62].
Collectively, many of these advances frame quantization as an optimization problem. Indeed, AdaRound’s method explicitly casts weight quantization as a QUBO instance [9]. These motivates our approach, which leverages quantum optimization to explore discrete quantization configurations. It’s important to note that these classical methods still operate within the standard non-convex training landscape. We aim to further reduce the quantization error beyond the reach of conventional techniques.
Convex formulation of neural network training.
The training of deep neural networks presents a formidable non-convex optimization challenge, limiting theoretical guarantees and interpretability [73]. A significant line of research has therefore sought to reformulate this problem within the framework of convex optimization, aiming to provide pathways to certified global optimality. For ReLU activation, Sahiner and Pilanci et al. [74, 73] employs convex duality to construct a semi-infinite program whose constraints are parameterized by the discrete activation sign patterns of the ReLU neurons. While exact, this formulation’s complexity scales exponentially with the rank of the data matrix, restricting its direct application to low-dimensional settings [74]. A second, distinct approach involves lifting the problem into a higher-dimensional matrix space, which circumvents the need for sign pattern enumeration [73].
Appendix B Error Bound of Piecewise Linear Approximation in FNN
Lemma 3
$n=O\left(\frac{1}{\sqrt{\epsilon}}\right).$ (87)
This result and analysis provide insight into the trade-offs between model efficiency and approximation accuracy, particularly relevant in the context of quantum-enhanced neural network training.